OpenStudy (anonymous):

Find the exact real value of arccos(1/2)

OpenStudy (turingtest):

have you tried it yet? it's just like the one we did

OpenStudy (anonymous):

cos=adj/hyp

OpenStudy (anonymous):

cos=1/2

OpenStudy (anonymous):

but i cant figure out the rest. im trying to see how arccos changes that or changes my answer

OpenStudy (turingtest):

actually sorry, this is just a special value

OpenStudy (turingtest):

not the same as the last :P my bad

OpenStudy (anonymous):

i thought so and plus the options dont match what i came up with.

OpenStudy (turingtest):

like I said in the last post, what is the definition of the arccosine?

OpenStudy (turingtest):

t=arccos(x) means what?

OpenStudy (anonymous):

arccos(X)= cos(x)

OpenStudy (turingtest):

no, the two functions are not the same... they are inverses

OpenStudy (turingtest):

arccos(x)=t implies that cos(t)=x

OpenStudy (turingtest):

so that means t is what kind of value?

OpenStudy (anonymous):

oops. i dont know what t is, im confused.

OpenStudy (anonymous):

t is for the triangle?

OpenStudy (turingtest):

normally when we have \[\cos\theta=x\] what does \(\theta\) represent?

OpenStudy (anonymous):

the theta is equal to the x or its just the value thats there

OpenStudy (turingtest):

theta is *not* the x, x is the sine of theta theta normally represents an *angle*, right?

OpenStudy (anonymous):

yes that makes more since

OpenStudy (turingtest):

so if\[\sin\theta=x\]and\[\sin^{-1}x=\theta\]then inverse trig functions are like asking "at what angle theta is the sine of that angle x?"

OpenStudy (turingtest):

so in your case the question becomes "the cosine of what angle(s) is/are 1/2 ?"

OpenStudy (anonymous):

i see so we have the hyp and the opp sides and we need the angle

OpenStudy (turingtest):

no, in this case this should be a value you have memorized the answer choices include pi unless they are in degrees, correct?

OpenStudy (anonymous):

i think it is pi/6

OpenStudy (turingtest):

so you are saying that cos(pi/6)=1/2 ?

OpenStudy (anonymous):

the options are: pi/3 2pi/3 -pi/6 pi/6

OpenStudy (anonymous):

no its pi/3

OpenStudy (turingtest):

correct :)

OpenStudy (anonymous):

thanks